# Data-Constrained Synthesis of Training Data for De-Identification

### v2

@@ 6 Conclusions
Data synthesis is an attractive tool for dealing with data scarcity and privacy risks. However, synthesis itself can be challenging when access to data is constrained. Through extensive ablation studies – validated on models and data in two different languages – our experiments show that not all parts of the synthesis process are equally sensitive to these resource constraints. Domain adapting the models to create high-utility clinical text did not require using all of the data. The experiments show that using between 25% and 50% of the data can be enough for domain adaptation, at least for these datasets. Furthermore, the experiments also show that using smaller generative LLMs does not necessarily incur a big loss of utility. Instead, our experiments show that the most important factor is the data available to create machine annotations.
These findings are a valuable contribution to the clinical NLP and privacy communities, where sharing real data is often impractical or impossible due to legal constraints. These constraints, while necessary from a privacy perspective, also hinder collaboration. While our studies cannot fully ascertain the safety of the synthetic corpora, the results indicate that they are less sensitive than the original data. When combined with other safety measures, such as de-identification and secure storage, synthetic data can serve as a basis for collaboration across institutional boundaries.
@@ 7 Limitations
Another limitation of the study is the reliance on n-gram-based metrics for estimating privacy risks. This is a common practice Ive et al. (2020); Hiebel et al. (2023) and can detect when data are being reproduced verbatim in the synthetic corpora. On the other hand, n-grams vary greatly in how sensitive they are. We try to address this with our n-gram recall metric that takes PII into account. However, we make limited claims about the privacy of the data and instead direct our focus towards their utility.
In Section 4.4, we find that smaller versions of the generative models could generate data of near-equal utility as their larger counterparts. This was especially clear when generating MEDDOCAN data. An interesting continuation would have been to fine-tune an even smaller LLM using MEDDOCAN data. Unfortunately, there is no smaller FLOR model than the 1.3 billion version that we use. GPT-SW3 is available in smaller versions, but proceeding to a monolingual analysis would lower the validity of the results and fell outside the scope of this study. Future work could try similar experiments with languages for which smaller models exist.
@@ 8 Ethical statement
This work was conducted under ethical permission no. 2019-05679 granted by Swedish Ethical Review Authority. MEDDOCAN is a publicly available dataset, where the PII in the documents are unrelated to the original patients. SEPR PHI is available on request and is a manually pseudonymized corpus where all identified PII have been replaced with surrogate values. Because of this, the privacy risks of the experiments in this paper are very small. Regardless, the experiments have been carried out in a computational environment in which only the authors and system administrators have had access to the data. Our experiments are also in accordance with the intended uses of the datasets.
The experiments conducted in this paper required considerable amounts of computational resources. We estimate that creating and evaluating the data and models for our experiments took approximately 130 GPU hours [^3] The calculations are available in Appendix C. Luckily, the computational infrastructure on which the experiments ran is located in a country where virtually all energy comes from sustainable sources. Nevertheless, the electricity expended when conducting these experiments could have been directed toward alternative purposes.
On the other hand, our results indicate that high-utility synthetic corpora can be created using small-scale data and without relying on the very largest LLMs. These results can be particularly helpful for researchers working in resource-constrained environments. This includes not only researchers in, e.g., the clinical domain, but also those working with low-resource languages. These parts of the NLP community are often under-served as increasing focus is placed on terabyte-scale datasets and LLMs with unwieldy parameter sizes.
There are other strong ethical benefits of conducting this study. Training models using synthetic corpora is safer than using real data, but is no panacea. On the other hand, no currently existing technique for privacy preservation is sufficient when used in isolation. For example, NER-based automatic de-identification covers only a subset of PII Pilán et al. (2022), and differentially private learning for NLP is difficult to implement properly Brown et al. (2022); Miranda et al. (2024) and is often inefficient when done so Igamberdiev et al. (2024). Synthetic data generation will likely be an important ingredient in many domains to overcome privacy issues.
There is a risk that other researchers over-interpret our results and use them to justify irresponsible uses of synthetic data. Our focus on constraining the amounts of data used hopefully mediates some of these potential risks. Furthermore, we have been clear in the limitations of our results and clearly defined the scope of our experiments.
@@ Appendix A Hyperparameters and software versions
Both GPT-SW3 and FLOR were fine-tuned using the hyperparameters in Table 8. These were selected based on examples from prior works and through small-scale experiments. SweDeClin-BERT and FLOR were also trained using the same hyperparameters, listed in Table 9. These parameters were also selected based on examples from prior works. Due to the many different factors in our synthesis process, a full search of the hyperparameter space was not feasible. The hyperparameters used in the experiments proved sufficiently optimized to obtain high-utility models and clear results.
@@ Appendix C Computational requirements
The experiments in this paper consumed many GPU hours due to the large number of configurations required for the ablation study. The GPUs were provided by National Academic Infrastructure for Supercomputing in Sweden. Unfortunately, the environment offers no straightforward way of computing the GPU hours. In this appendix, we estimate the GPU hours required to run the experiments based on data from the logs.

### v3

@@ 6 Conclusions
Data synthesis is an attractive tool for dealing with data scarcity and privacy risks. However, synthesis itself can be challenging when access to data is constrained. Through extensive ablation studies – validated on models and data in two different languages – our experiments show that not all parts of the synthesis process are equally sensitive to these resource constraints. Domain adapting the models to create high-utility clinical text did not require using all of the data. The experiments show that using between 25% and 50% of the data can be enough for domain adaptation, at least for the studied datasets. Furthermore, the experiments also show that using smaller generative LLMs does not necessarily incur a big loss of utility. Instead, our experiments show that the most important factor is the data available to create machine annotations.
These findings are a valuable contribution to the clinical NLP and privacy communities, where sharing real data is often impractical or impossible due to legal constraints. These constraints, while necessary from a privacy perspective, also hinder collaboration. While our studies cannot fully ascertain the privacy of the synthetic corpora, the results indicate that they are less sensitive than the original data. When combined with other safety measures, such as de-identification and secure storage, synthetic data can serve as a basis for collaboration across institutional boundaries.
@@ 7 Limitations
Another limitation of the study is the reliance on n-gram-based metrics for estimating privacy risks. This is a common practice Ive et al. (2020); Hiebel et al. (2023) and can detect when data are being reproduced verbatim in the synthetic corpora. On the other hand, n-grams vary greatly in how sensitive they are. We try to address this with our n-gram recall metric that takes PII into account. However, we make limited claims about the privacy of the data and instead focus on their utility.
In Section 4.4, we find that smaller versions of the generative models could generate data of near-equal utility as their larger counterparts. This was especially clear when generating MEDDOCAN data. An interesting continuation would have been to fine-tune an even smaller LLM using MEDDOCAN data. Unfortunately, the 1.3 billion version of FLOR that we use is the smallest one. GPT-SW3 is available in smaller versions, but proceeding to a monolingual analysis would lower the validity of the results and fell outside the scope of this study. Future work could try similar experiments with languages for which smaller models exist.
@@ 8 Ethical Statement
This work was conducted under ethical permission no. 2019-05679 granted by the Swedish Ethical Review Authority. MEDDOCAN is a publicly available dataset, where the PII in the documents are unrelated to the original patients. SEPR PHI is available on request and is a manually pseudonymized corpus where all identified PII have been replaced with surrogate values. Because of this, the privacy risks of the experiments in this paper are very small. Regardless, the experiments have been carried out in a computational environment in which only the authors and system administrators have had access to the data. Our experiments are also in accordance with the intended uses of the datasets.
The experiments conducted in this paper required considerable amounts of computational resources. We estimate that creating and evaluating the data and models for our experiments took approximately 130 GPU hours [^4] The calculations are available in Appendix C. Luckily, the experiments were run in Sweden – where virtually all energy comes from sustainable sources. Nevertheless, the electricity expended when conducting these experiments could have been used for other purposes.
On the other hand, our results indicate that high-utility synthetic corpora can be created using small-scale data and without relying on the very largest LLMs. These results can be particularly helpful for researchers working in resource-constrained environments. This includes not only researchers in, e.g., the clinical domain, but also those working with low-resource languages. These parts of the NLP community are often under-served as increasing focus is placed on terabyte-scale datasets and LLMs with unwieldy amounts of parameters.
Even though synthetic data are safer than sensitive data, there is a risk that other researchers over-interpret our results and use them to justify irresponsible uses of synthetic data. Our focus on constraining the amounts of data used hopefully mediates some of these potential risks. Furthermore, we have thoroughly described the limitations of our results and the scope of our experiments.
@@ Appendix A Hyperparameters and Software Versions
Both GPT-SW3 and FLOR were fine-tuned using the hyperparameters in Table 8. These were selected based on examples from prior works and through small-scale experiments. SweDeClin-BERT and roberta-base-bne were also trained using the same hyperparameters, listed in Table 9. These parameters were also selected based on examples from prior works. If the encoder models got stalled in the local minimum of only predicting the zero class, then training was restarted using the same data and hyperparameters. Due to the many different factors in our synthesis process, a full search of the hyperparameter space was not feasible. The hyperparameters used in the experiments proved sufficiently optimized to obtain high-utility models and clear results.
@@ Appendix C Computational Requirements
The experiments in this paper consumed many GPU hours due to the large number of configurations required for the ablation study. The GPUs were provided by the National Academic Infrastructure for Supercomputing in Sweden. Unfortunately, the environment offers no straightforward way of computing the GPU hours. In this appendix, we estimate the GPU hours required to run the experiments based on data from the logs.